### v1

@@ 5.4 Ablation experiment
A comprehensive ablation experiment is designed to validate the contribution of each part of our proposed model AHMDF. We remove module (a), (b), and (c), price information, temperature information, and variable selection networks, respectively. Specifically, we also replace Gumbel-Softmax with Softmax in the gate temporal attention module to demonstrate its usefulness. In Table 3, we present the results of ablation experiments, which are conducted in 4 different intervals (i.e., 15, 30, 45, and 60 min). The full model, AHMDF, improves about 10.04% in RMSE, 15.85% in RAE, and 15.85% in MAE on average. In the 60-minute prediction, the proposed model AHMDF is weakly worse than “# Gumbel-Softmax” in $\rm R^{2}$, but better in the other three metrics. We still argue that the ablation experiment can be passed because its overall performance is not bad. Different modules and information contribute differently to improving prediction accuracy. First, Module (b) (i.e. adjacency graph fusion module) contributes the most to the prediction, demonstrating the importance of learning adjacency features. Second, Module (c) (i.e. temporal feature fusion module) is helpful in improving long-time prediction accuracy, and removing Module (c) sharply deteriorates the results as the prediction interval increases. Besides, time-series information, such as prices and temperatures, both contribute to the accuracy of long-term prediction. Third, module (a) enables the model to effectively capture the dynamic changes of the areas with same land use, thus improving prediction accuracy. Fourth, the variable selection network helps to determine the weights among the temporal variables, and the gated temporal attention allows a strong focus on important temporal features, both of which provide a non-negligible boost.
@@ 5.5 Spatial correlation analysis
To illustrate the effectiveness of our geo-based clustering method, we show the true correlation of 247 areas’ occupancies and the connectivity relationship obtained by our clustering method in Fig. 6. Fig. 6(a) illustrates the areas that show a strong positive correlation. We consider a relatively strong correlation when the Pearson correlation coefficient between the two areas is greater than or equal to 0.4. Fig. 6(b) illustrates the relations connected by hyperedges in our hypergraph. A black square indicates that the two areas are connected by a hyperedge, and conversely white one indicates that they are not connected. The two images are similar in many places, which shows that our clustering method can effectively capture spatial data correlation. To give concrete examples, the red box shows some residential areas in Luohu District, Shenzhen, and the purple box shows some commercial areas in Futian District, Shenzhen. Areas with the same land use type have similar patterns of changing charging demand as shown in Fig. 6(a). Corresponding to that, our approach makes good use of POIs to cluster areas with similar land use attributes and construct hyperedge connectivity relationships.
@@ 5.6 Hyperparameter analysis
In this subsection, we discuss the impact of value choices of number of adaptive areas clusters $C$, number of encoder components $N_{E}$ and temperature coefficient of Gumbel-Softmax $\varpi$ on the results. First, as shown in Fig. 7, despite $C$ being set differently, the area adaptive clustering roughly matches reality. The southern urban built-up area, developing north-western, north-eastern and central parts, and forests, reservoirs and city parks are clustered separately. However, when $C$ is too small (e.g., $C=6$), the method has difficulty in distinguishing between urban sub-centers and suburban tourist areas. Besides, when $C$ is too large, the method tends to cut the city into too many patches, e.g., as many as six types of functional areas in the southern CBD area when $C=14$, which is unfavorable for subsequent hypergraph studies. Second, we determine the appropriate hyperparameters by searching experiment, as shown in Fig. 8. Our experiment results demonstrate the robustness of the proposed model AHMDF, as it is insensitive to the change of hyperparameters. $N_{E}$ affects the depth of the network. Networks that are too shallow may have difficulty capturing complex nonlinear relationships, while networks that are too deep tend to cause overfitting and waste of computational resources. $\varpi$ determines the number of temporal features that Gumbel-Softmax focuses on. A too small $\varpi$ makes the model tend to one-hot attention, causing insufficient feature attention, while too large $\varpi$ causes Gumbel-Softmax to converge to softmax, causing fuzzy attention that we wish to avoid from the beginning. Therefore, we choose the hyperparameter setting that is most favorable for performance, as shown in Subsection 5.2.
In summary, we propose AHMDF, a deep learning model that incorporates dynamic and static influences, for accurate citywide EV charging demand prediction. The experimental results empirically show that AHMDF achieves state-of-the-art, with an average improvement of 9.75% in three metrics and achieves the best fit in $\rm R^{2}$. Specifically, our method performs well on relatively long time prediction. The ablation experiments demonstrate the joint contribution of each module, and the full model achieves a 12.26% improvement over the ablation models. We also explore the role of dynamic factors in improving the prediction performance of different city areas. Furthermore, we experimentally explored the impact of hyperparameters.
@@ 6 Conclusion
Accurate EV charging demand prediction is an important foundation for smart grids and transportation systems, helping to optimize the grid and guide the demand. How to incorporate multi-source heterogeneous external data into the model so that it can have a positive impact on improving the accuracy of EV charging demand prediction is a current research hotspot. To address the remaining research challenges, we propose an attention-based approach named AHMDF that incorporates attentive hypergraph neural network for classified areas groups, adjacency graph attention and an improved Transformer encoder. AHMDF (1) outperforms 10 representative methods by 9.75% on average in three metrics and achieves the best fit in $\rm R^{2}$; (2) outperforms 7 ablation models comprehensively, demonstrating the enhancement brought by each introduced multi-source data or module. Moreover, we analyze the effect of hyperparameters and determined the preferred values in our case.
Furthermore, our research provides some perspectives on the management of smart energy and electric vehicle charging. (1) Achieving accurate charging demand prediction requires not only a focus on the charging data itself, but also on the external factors that influence EV charging choices. Therefore, fusion of information from multiple sources at the government or manager level will be of significant help in collaborative and refined management. (2) Incorporating external static and dynamic influences in the model may be of great help in achieving accurate long term prediction. (3) Among the dynamic influences, price information contributes significantly to the accurate prediction of residential areas, while temperature information contributes to that of suburban and tourist areas. (4) Associating charging facilities with similar properties in the surrounding helps in accurate node information propagation and collaborative area management.

### v2

@@ 4.4 Ablation experiment
A comprehensive ablation experiment is designed to validate the contribution of each part of our proposed model CityEVCP. We remove module (a), (b), and (c), price information, temperature information, and variable selection networks, respectively. Specifically, we also replace Gumbel-Softmax with Softmax in the gate temporal attention module to demonstrate its usefulness. In Table 3, we present the results of ablation experiments, which are conducted in 4 different intervals (i.e., 15, 30, 45, and 60 min). The full model, CityEVCP, improves about 10.04% in RMSE, 15.85% in RAE, and 15.85% in MAE on average. In the 60-minute prediction, the proposed model CityEVCP is weakly worse than “# Gumbel-Softmax” in $\rm R^{2}$, but better in the other three metrics. We still argue that the ablation experiment can be passed because its overall performance is not bad. Different modules and information contribute differently to improving prediction accuracy. First, Module (b) (i.e. adjacency graph fusion module) contributes the most to the prediction, demonstrating the importance of learning adjacency features. Second, Module (c) (i.e. temporal feature fusion module) is helpful in improving long-time prediction accuracy, and removing Module (c) sharply deteriorates the results as the prediction interval increases. Besides, time-series information, such as prices and temperatures, both contribute to the accuracy of long-term prediction. Third, module (a) enables the model to effectively capture the dynamic changes of the areas with same land use, thus improving prediction accuracy. Fourth, the variable selection network helps to determine the weights among the temporal variables, and the gated temporal attention allows a strong focus on important temporal features, both of which provide a non-negligible boost.
@@ 4.5 Spatial correlation analysis
To illustrate the effectiveness of our geo-based clustering method, we show the true correlation of 247 areas’ occupancies and the connectivity relationship obtained by our clustering method in Fig. 7. Fig. 7(a) illustrates the areas that show a strong positive correlation. We consider a relatively strong correlation when the Pearson correlation coefficient between the occupancy data of two areas is greater than or equal to 0.4. Fig. 7(b) illustrates the relations connected by hyperedges in our hypergraph. A black square indicates that the two areas are connected by a hyperedge, and conversely white one indicates that they are not connected. The two images are similar in many places, which shows that our clustering method can effectively capture spatial data correlation. To give concrete examples, the red box shows some residential areas in Luohu District, Shenzhen, and the purple box shows some commercial areas in Futian District, Shenzhen. Areas with the same land use type have similar patterns of changing charging demand as shown in Fig. 7(a). Corresponding to that, our approach makes good use of POIs to cluster areas with similar land use attributes and construct hyperedge connectivity relationships.
@@ 4.6 Hyperparameter analysis
In this subsection, we discuss the impact of hyperparameters (i.e., number of adaptive areas clusters $C$, number of encoder components $N_{E}$ and temperature coefficient of Gumbel-Softmax $\varpi$). First, as shown in Fig. 8, despite $C$ being set differently, the area adaptive clustering roughly matches reality. The southern urban built-up area, developing north-western, north-eastern and central parts, and forests, reservoirs and city parks are clustered separately. However, when $C$ is too small (e.g., $C=6$), the method has difficulty in distinguishing between urban sub-centers and suburban tourist areas. Besides, when $C$ is too large, the method tends to cut the city into too many patches, e.g., as many as six types of functional areas in the southern CBD area when $C=14$, which is unfavorable for subsequent hypergraph studies. Second, we determine the appropriate hyperparameters by searching experiment, as shown in Fig. 9. Our experiment results demonstrate the robustness of the proposed model CityEVCP, as it is insensitive to the change of hyperparameters. $N_{E}$ affects the depth of the network. Networks that are too shallow may have difficulty capturing complex nonlinear relationships, while networks that are too deep tend to cause overfitting and waste of computational resources. $\varpi$ determines the number of temporal features that Gumbel-Softmax focuses on. A too small $\varpi$ makes the model tend to one-hot attention, causing insufficient feature attention, while too large $\varpi$ causes Gumbel-Softmax to converge to softmax, causing fuzzy attention that we wish to avoid from the beginning. Therefore, we choose the hyperparameter setting that is most favorable for performance, as shown in Subsection 4.2.
In summary, we propose CityEVCP, a deep learning model that incorporates dynamic and urban region influences, to achieve accurate citywide EV charging demand prediction. The experimental results empirically show that CityEVCP achieves state-of-the-art, with an average improvement of 10.47% in three metrics and achieves the best fit in $\rm R^{2}$. Specifically, our method performs well on relatively long time prediction. The ablation experiments demonstrate the joint contribution of each module, and the full model achieves a 13.91% improvement over the ablation models. We also explore the role of dynamic factors in improving the prediction performance of different city areas and illustrates the correct understanding of the proposed model. Furthermore, we experimentally explored the impact of hyperparameters.
@@ 5 Conclusion
Accurate EV charging demand prediction is an important foundation for smart grids and transportation systems, helping to optimize the grid and guide the demand. How to incorporate multi-source heterogeneous external data into the model so that it can have a positive impact on improving the accuracy of EV charging demand prediction is a current research hotspot. To address the remaining research challenges, we propose an data-driven approach named CityEVCP that incorporates attentive hypergraph neural network for classified areas groups, adjacency graph attention, variable selection network and an improved Transformer encoder. CityEVCP (1) outperforms 10 representative methods by 10.47% on average in three metrics and achieves the best fit in $\rm R^{2}$; (2) outperforms 7 ablation models comprehensively, demonstrating the enhancement brought by each introduced multi-source data or module. In our case study, we analyze the role of charging prices and temperatures for different areas of the city, validating the model’s ability to correctly understand the dynamic auxiliary information. Moreover, we analyze the impact of hyperparameters and determined the preferred values in our case.
Furthermore, our research provides some perspectives on the management of smart energy and electric vehicle charging. First, achieving accurate charging demand prediction requires not only a focus on the charging data itself, but also on the external factors that influence EV charging choices. Therefore, information fusion from multiple sources at the government or manager level will be of significant help in collaborative and refined management. Second, among the dynamic influences, price information contributes significantly to the accurate prediction of residential areas, while temperature information contributes to that of suburban and tourist areas. Third, associating charging facilities with similar properties in the surrounding helps in accurate node information propagation and collaborative area management.